A positive number when divided by 35 gives the remainder 33. What will the remainder be if this number is divided by 7?
Answer by richwmiller(17219)

x/35=z+33/35
x=35*(z+33/35)
x=35z+33
x/7=35z/7+33/7
5z+33/7=(5z+4)+5/7
The remainder will be 5